OpenStudy (anonymous):

There’s a game that I really enjoy playing where the object is to knock down various pigs with birds. The cool thing about the game is that, in order to master it, you must have a solid understanding of parabolas. To knock down a certain pig, I think that my quadratic equation would be y= 2x2 + 4x + 3. Are there one or two possible solutions to this problem? How do you know? The flights all will begin and end on the x axis. Assuming that this is a flight, where does the bird begin her flight (assuming left to right flight)? (Round to the nearest hundredth, if needed.)

OpenStudy (precal):

angry birds is your game, but you can use other functions other than a parabola. horizontal line - straight across vertical line-up how about the grey little bird that breaks up into 3 birds

OpenStudy (precal):

now if you are solving that particular quadratic, just use the quadratic formula

OpenStudy (anonymous):

ok, I know nothing about the game angry birds, so I am totally lost with that part. This week we are studying parabola so I think that is the way that I shoud figure this out or am i wrong on that?

OpenStudy (precal):

so you made up the equation?

OpenStudy (anonymous):

I am trying as hard as I can to get this right, if I don't I'm going to flunk my mat/117 class and I don't want to fail yes the equation is y = - b/2a, which is -(4)/2(2) = -1 2(-1)^2 + (-4) + 3 = 3

OpenStudy (precal):

that is the vertex? Is that what you are looking for?

OpenStudy (precal):

this is a max point

OpenStudy (anonymous):

I found the vertex and the line of symmetry, now I need to find out the flight, left to right, of the bird, and I'm lost on that one

OpenStudy (precal):

use the quadratic formula

OpenStudy (anonymous):

plus i need to figure out if there is one or two solutions, I'm thinking that there is only one, am I right?

OpenStudy (anonymous):

you meant instead of the formula I said before?

OpenStudy (precal):

well solutions mean how many times it crosses the x axis

OpenStudy (precal):

 I would call it two

OpenStudy (precal):